Interest rate​ (with changing​ years). Keiko is looking at the following investment choices and wants to know what annual rate of return each choice produces. a.  Invest ​$450.00 and receive ​$704.33 in 8 years. b.  Invest ​$3,800.00 and receive ​$10,820.09 in 15 years. c.  Invest ​$31,234.32 and receive ​$140,000.00 in 20 years. d.  Invest ​$33 ,378.89 and receive ​$1,000,000.00 in 45 years.
Answer #1

a

Future value = present value*(1+ rate)^time
704.33 = 450*(1+Interest rate/100)^8
Interest rate % = 5.76

b

Future value = present value*(1+ rate)^time
10820.09 = 3800*(1+Interest rate/100)^15
Interest rate % = 7.23

c

Future value = present value*(1+ rate)^time
140000 = 31234.32*(1+Interest rate/100)^20
Interest rate % = 7.79

d

Future value = present value*(1+ rate)^time
1000000 = 33378.89*(1+Interest rate/100)^45
Interest rate % = 7.85

Choose D